# All-in-One: Heterogeneous Interaction Modeling for Cold-Start Rating Prediction

### v2

@@ VI-B Comparison Results
We investigate the overall performance of HIRE in 3 cold-start scenarios, i.e., user cold-start (UC), item cold-start (IC), and user & item cold-start (U&I C), in comparison with the 8 baseline approaches. Table III, IV and V list the test results on three datasets, respectively, where the first-best (yellow) and the second-best (underlined) scores are highlighted. In general, HIRE achieves the best test accuracy in most cases. The Precision, NDCG and MAP of HIRE succeed all the baselines 0.22, 0.30, 0.22 on average, respectively. The superiority of HIRE is reflected in improving NDCG while keeping high Precision and MAP.
In contrast to neural CF-based approaches, NeuMF, Wide&Deep, DeepFM and AFN, HIRE outperforms them on all the datasets significantly. The CF-based approaches only model a single type of user-item interaction via observed ratings, which is difficult to generalize to new users/items with only a few interactions. This reflects modeling heterogeneous interactions is beneficial for cold-start prediction.
@@
The meta-learning approaches, MAMO, TaNP and MeLU, outperform the CF-based approaches by a large margin. The performance of these three approaches dominates the second best result, and is competitive to our HIRE in some cases. The meta-learning approaches focus on exploiting parameter sharing and adaption to address cold-start problem. Their complicated adaption module and learning strategy may lead to a difficult learning procedure. HIRE deals with the cold-start problem in a different way that relies on the interaction learned from the data. It concentrates on using a relatively simple neural network framework and learning algorithm to model heterogeneous interactions in a data-driven fashion.
@@
On $\mathsf{MovieLens}$-$\mathsf{1M}$ (Table III), we compared HIN-based baseline GraphHINGE and MetaHIN. In the user cold-start scenario, GraphHINGE gains better performance than CF-based approaches, indicating that metapath-guided neighborhood can capture the complex and high-order semantics in the HIN. In item cold-start and user & item cold-start scenarios, MetaHIN consistently outperforms GraphHINGE, as indicated by its higher NDCG scores. However, these two approaches fail to outperform HIRE and TaNP. It validates again that manually defined heterogeneous patterns may not contribute to estimating preference, while the learned heterogeneous interactions by HIRE tend to be more reliable.
We compare the social recommendation baseline, GraphRec, on $\mathsf{Douban}$ (Table V). GraphRec performs better in the user cold-start scenario, and can even surpasses HIRE. GraphRec uses Graph Neural Network to integrate the rating interactions and social relations, which generates better user representations. However, we find that GraphRec is less effective in scenarios with cold items, where the social relations between users are less helpful for cold items.
Figure 6: Total Test Time (second)
VI-C Efficiency
@@
TABLE VI: Ablation Study for Different Attention Layers on $\mathsf{MovieLens}$-$\mathsf{1M}$
@@
wo/ Item & Attribute 0.4465 0.7858 0.3232 0.4392 0.7600 0.3177 0.4663 0.7700 0.3440
wo/ User & Attribute 0.6552 0.8926 0.5838 0.5268 0.8174 0.4301 0.5227 0.8138 0.4239
wo/ User & Item 0.6752 0.8986 0.6040 0.5163 0.8128 0.4202 0.5067 0.8079 0.4073
wo/ User 0.6590 0.8925 0.5885 0.5272 0.8116 0.4223 0.5239 0.8111 0.4213
wo/ Item 0.4461 0.7866 0.3238 0.4414 0.7610 0.3193 0.4687 0.7700 0.3447
wo/ Attribute 0.4477 0.7865 0.3242 0.4413 0.7611 0.3200 0.4671 0.7699 0.3442
full model 0.6787 0.9002 0.6097 0.5871 0.8475 0.4993 0.5848 0.8493 0.5008
@@ VI-E Ablation Study
Impacts of the three types of attention layers. We conduct an ablation study to investigate the effectiveness of the 3 types of attention layers, attention between users, items and attributes, in HIRE. Here, we train 6 model variants on $\mathsf{MovieLens}$-$\mathsf{1M}$, where single types or combinations of two types of layers are removed for the model. Table VI lists the test performance of the model variants in 3 cold-start scenarios. The full model with the 3 types of attention layers achieves the overall best performance, indicating that the heterogeneous interactions from 3 different perspectives collaboratively contribute to the rating prediction. In addition, we find that the attention between items (or attributes) plays a more important role than attention between users in rating prediction, by comparing model variants wo/ Item (or wo/ Attribute) with model variant wo/ User. For $\mathsf{MovieLens}$-$\mathsf{1M}$, we speculate that the similarity between the movies and the similarity between the proprieties of users and movies influences the rating to a larger degree. Meanwhile, another possible reason would be the interactions between users may not be reliable. That is because there is a counter-intuitive observation on that the model variant wo/ Item (or wo/ Attribute) performs worse than the model variant wo/ User & Item (or wo/ User & Attribute). And the model with only attention between users, i.e., wo/ Item & Attribute, leaves the worst performance among all the model variants. The attention between users may learn irrelevant interactions, where deploying this layer alone or combining it with another single type of attention layer incurs extra noise for the prediction. Furthermore, in our experiments on $\mathsf{Bookcrossing}$ and $\mathsf{Douban}$ datasets, the full model consistently outperforms other variants.
@@ VI-F Case Study
We conduct a case study to investigate whether HIM learns relevant and reliable interactions between users, items and attributes, and here we denote the three layers as MBU, MBI and MBA, respectively. We visualize the attention matrices for the prediction on $\mathsf{MovieLens}$-$\mathsf{1M}$ in Fig. 9. Here, the darker the cell, the higher the corresponding weight, and the stronger the implicit interaction learned by HIM. First, Fig. 9(a) shows the attention weights in MBU layer among 16 users for item $1,304$, which is an action and comedy movie, ’Butch Cassidy and the Sundance Kid’. In Fig. 9(a), as the highlighted red rectangle, user $1,447$ is deeply affected by user $4,926$ and $3,792$. We find that the three users share the same preference on item $1,304$. With the learned user interactions, HIRE predicts the ratings of the three users on item $1,304$ as $3.41$, $4.13$ and $4.20$, which is highly consistent to the ground-truth ratings, $3,4,4$, respectively. That indicates the interaction between users has a strong correlation to their ratings. Second, Fig. 9(b) shows the attention weights in MBI layer among 16 items for user $4,926$, who is a female student under 18 years old. Similarly, we observe that item $487$ is strongly affected by items $1,196$, $1,197$ and $1,198$. With the learned interaction between items, HIRE predicts the ratings of that user on item $487$ as $3.42$, and on items $1,196$, $1,197$ and $1,198$ as $3.88,3.95,3.86$, respectively. That is also consistent to the ground-truth ratings of $3$ for item $487$ and $4$ for the other three items. The results also demonstrate that the learned implicit interactions among items guide the model to predict similar ratings for interacted items. Finally, Fig. 9(c) and 9(d) present the attention weights among the 8 attributes in the MBA layer, for two user-item pairs, user 4,926 and item 1,304, user 1,051 and item 2,081, respectively. The former pair has a high rating of $4$. User $1051$ is a $18-24$ year old male technician/engineer and item $2081$ is a comedy ’Trees Lounge’. The user $1051$ gave a low rating $2$ to movie $2,081$. It is reasonable according to the occupation of user and genre of movie. As we can see, high rating pair has more attribute interactions. Actor attribute has interactions with occupation in Fig. 9(c). This means that the actor in the movie may be liked by people in a specific occupation. The latter one has less interactions between movie attributes and user attributes. It may because user has no interest to this movie. It is worth mentioning that the weight matrices are unsymmetrical due to the computation of attention (Eq. (2)), which indicates the interactions are single direction.
@@ VII Conclusion
In this paper, we propose a fully data-driven recommender system HIRE that can model heterogeneous interactions for cold-start rating prediction. For target cold user/items, we adopt a neighborhood-based sampling strategy to sample prediction context from the user-item bipartite rating graph, and explore reliable interactions in the context. Specifically, we devise a Heterogeneous Interaction Module to jointly model the interactions in three levels, i.e., users, items and attributes, respectively. Comprehensive experiments on three real-world datasets validate the effectiveness of our model. HIRE outperforms CF-based approaches, social recommendation approaches, HIN-based approaches and state-of-the-art approaches with higher precision, NDCG and MAP by 0.22, 0.30 and 0.22 on average.

### v3

@@ VI-B Comparison Results
We investigate the overall performance of HIRE in 3 cold-start scenarios, i.e., user cold-start (UC), item cold-start (IC), and user & item cold-start (U&I C), in comparison with the 12 baseline approaches. We also investigate whether HIRE can effectively handle regular cases by using data from warm states and we compare HIRE against several competitive baselines. This means that the test users and items may also appear in the training dataset and have a substantial history of interactions. The results are reported in Table II. Table III, IV and V list the test results on 3 datasets, respectively, where the first-best (yellow) and the second-best (underlined) scores are highlighted. In general, HIRE achieves the best test accuracy in most cases. The Precision, NDCG and MAP of HIRE succeed all the baselines 0.21, 0.29, 0.22 on average, respectively. The superiority of HIRE is reflected in improving NDCG while keeping high Precision and MAP.
In Table II, HIRE outperforms these baselines, demonstrating its capability in handling regular warm cases effectively. Furthermore, since in regular cases, the prediction context involves a substantial amount of historical interactions for the test users/items, HIRE achieves better performance in regular cases than in the cold-start scenarios in most cases.
@@
Popularity achieves a very competitive performance on $\mathsf{MovieLens}$-$\mathsf{1M}$ and $\mathsf{Douban}$, since the correlation between the number of interactions and the rating is positive in these datasets. TransGNN does not perform well in cold-start scenarios since it relies heavily on user-item interactions. In contrast to neural CF-based approaches, NeuMF, Wide&Deep, DeepFM and AFN, HIRE outperforms them on all the datasets significantly. The CF-based approaches only model a single type of user-item interaction via observed ratings, which is difficult to generalize to new users/items with only a few interactions. This reflects modeling heterogeneous interactions is beneficial for cold-start prediction. The meta-learning approaches, MAMO, TaNP and MeLU, outperform the CF-based approaches by a large margin. The performance of these 3 approaches dominates the second best result, and is competitive to our HIRE in some cases. The meta-learning approaches focus on exploiting parameter sharing and adaption to address cold-start problem. Their complicated adaption module and learning strategy may lead to a difficult learning procedure. HIRE deals with the cold-start problem in a different way that relies on the interaction learned from the data. It concentrates on using a relatively simple neural network framework and learning algorithm to model heterogeneous interactions in a data-driven fashion.
@@
On $\mathsf{MovieLens}$-$\mathsf{1M}$ (Table III), we compared HIN-based baseline GraphHINGE and MetaHIN. In the user cold-start scenario, GraphHINGE gains better performance than CF-based approaches, indicating that metapath-guided neighborhood can capture the complex and high-order semantics in the HIN. In item cold-start and user & item cold-start scenarios, MetaHIN consistently outperforms GraphHINGE, as indicated by its higher NDCG scores. However, these two approaches fail to outperform HIRE and TaNP. It validates again that manually defined heterogeneous patterns may not contribute to estimating preference, while the learned heterogeneous interactions by HIRE tend to be more reliable. We compare the social recommendation baseline, GraphRec, on $\mathsf{Douban}$ (Table V). GraphRec performs better in the user cold-start scenario, and can even surpasses HIRE. GraphRec uses Graph Neural Network to integrate the rating interactions and social relations, which generates better user representations. However, we find that GraphRec is less effective in scenarios with cold items, where the social relations between users are less helpful for cold items.
Figure 6: Total Test Time (in second)
@@
TABLE VI: Ablation Study for Different Attention Layers on $\mathsf{MovieLens}$-$\mathsf{1M}$ (%)
@@
wo/ Item & Attribute 44.65 78.58 32.32 43.92 76.00 31.77 46.63 77.00 34.40
wo/ User & Attribute 65.52 89.26 58.38 52.68 81.74 43.01 52.27 81.38 42.39
wo/ User & Item 67.52 89.86 60.40 51.63 81.28 42.02 50.67 80.79 40.73
wo/ User 65.90 89.25 58.85 52.72 81.16 42.23 52.39 81.11 42.13
wo/ Item 44.61 78.66 32.38 44.14 76.10 31.93 46.87 77.00 34.47
wo/ A 44.77 78.65 32.42 44.13 76.11 32.00 46.71 76.99 34.42
wo/ Attribute 69.99 91.69 64.54 59.89 86.40 53.04 60.30 86.93 53.62
VI-C Efficiency
@@ VI-E Ablation Study
Impacts of the three types of attention layers. We conduct an ablation study to investigate the effectiveness of the three types of attention layers, attention between users, items and attributes, in HIRE. Here, we train 6 model variants on $\mathsf{MovieLens}$-$\mathsf{1M}$, where single types or combinations of two types of layers are removed for the model. Table VI lists the test performance of the model variants in 3 cold-start scenarios. The full model with the three types of attention layers achieves the overall best performance, indicating that the heterogeneous interactions from three different perspectives collaboratively contribute to the rating prediction. In addition, we find that the attention between items (or attributes) plays a more important role than attention between users in rating prediction, by comparing model variants wo/ Item (or wo/ Attribute) with model variant wo/ User. For $\mathsf{MovieLens}$-$\mathsf{1M}$, we speculate that the similarity between the movies and the similarity between the proprieties of users and movies influences the rating to a larger degree. Meanwhile, another possible reason would be the interactions between users may not be reliable. That is because there is a counter-intuitive observation on that the model variant wo/ Item (or wo/ Attribute) performs worse than the model variant wo/ User & Item (or wo/ User & Attribute). And the model with only attention between users, i.e., wo/ Item & Attribute, leaves the worst performance among all the model variants. The attention between users may learn irrelevant interactions, where deploying this layer alone or combining it with another single type of attention layer incurs extra noise for the prediction. Furthermore, in our experiments on $\mathsf{Bookcrossing}$ and $\mathsf{Douban}$ datasets, the full model consistently outperforms other variants.
@@ VI-F Case Study
We conduct a case study to investigate whether HIM learns relevant and reliable interactions between users by the 3 layers MBU, MBI and MBA. We visualize the attention matrices for the prediction on $\mathsf{MovieLens}$-$\mathsf{1M}$ in Fig. 9. Here, the darker the cell, the higher the corresponding weight, and the stronger the implicit interaction learned by HIM. First, Fig. 9(a) shows the attention weights in MBU layer among 16 users for item $1,304$, which is an action and comedy movie, ‘Butch Cassidy and the Sundance Kid’. In Fig. 9(a), as the highlighted red rectangle, user $1,447$ is deeply affected by user $4,926$ and $3,792$. We find that the three users share the same preference on item $1,304$. With the learned user interactions, HIRE predicts the ratings of the three users on item $1,304$ as $3.41$, $4.13$ and $4.20$, which is highly consistent to the ground-truth ratings, $3,4,4$, respectively. That indicates the interaction between users has a strong correlation to their ratings. Second, Fig. 9(b) shows the attention weights in MBI layer among 16 items for user $4,926$, who is a female student under 18 years old. Similarly, we observe that item $487$ is strongly affected by items $1,196$, $1,197$ and $1,198$. With the learned interaction between items, HIRE predicts the ratings of that user on item $487$ as $3.42$, and on items $1,196$, $1,197$ and $1,198$ as $3.88,3.95,3.86$, respectively. That is also consistent to the ground-truth ratings of $3$ for item $487$ and $4$ for the other three items. The results also demonstrate that the learned implicit interactions among items guide the model to predict similar ratings for interacted items. Finally, Fig. 9(c) and 9(d) present the attention weights among the 8 attributes in the MBA layer, for two user-item pairs, user 4,926 and item 1,304, user 1,051 and item 2,081, respectively. The former pair has a high rating of $4$. User $1051$ is a $18-24$ year old male technician/engineer and item $2081$ is a comedy ‘Trees Lounge’. The user $1051$ gave a low rating $2$ to movie $2,081$. It is reasonable according to the occupation of user and genre of movie. As we can see, high rating pair has more attribute interactions. Actor attribute has interactions with occupation in Fig. 9(c). This means that the actor in the movie may be liked by people in a specific occupation. The latter one has less interactions between movie attributes and user attributes. It may because user has no interest to this movie. It is worth mentioning that the weight matrices are unsymmetrical due to the computation of attention (Eq. (2)), which indicates the interactions are single direction.
@@ VII Conclusion
In this paper, we propose a fully data-driven recommender system HIRE that can model heterogeneous interactions for cold-start rating prediction. For target cold user/items, we adopt a neighborhood-based sampling strategy to sample prediction context from the user-item bipartite rating graph, and explore reliable interactions in the context. Specifically, we devise a Heterogeneous Interaction Module to jointly model the interactions in three levels, i.e., users, items and attributes, respectively. Comprehensive experiments on 3 real-world datasets validate the effectiveness of our model. HIRE outperforms baselines with higher precision, NDCG and MAP by 0.21, 0.29 and 0.22 on average. The source code is publicly available at https://github.com/FangShuheng/HIRE.
Acknowledgments
This work was supported by the Research Grants Council of Hong Kong, China, No. 14203618, No. 14202919 and No. 14205520. Kangfei Zhao is supported by National Key Research and Development Plan, No. 2023YFF0725101.